OpenStudy (anonymous):

Using substitution to solve the system x = 2y + 3 and 4x - 5y = 9, which of the following could be the first step? A. x - 2y = 3 4x - 5y = 9 B. 4(2y + 3) - 5y = 9 C. x + 2y = 4x - 5y D. 5x - 3y = 12 E. - 3x + 7y = -6

OpenStudy (callisto):

put x=2y+3 into 4x - 5y = 9 4(2y+3) - 5y = 9

OpenStudy (callisto):

i hope you understand what the meaning of substitution is..

OpenStudy (anonymous):

substitution is when u place the value of something into an equation. In this case you have x equals 2y + 3 so u can put that into the equation 4x - 5y = 9 to get 4(2y+3) - 5y = 9